OpenStudy (anonymous):

Solve this system of equations using any method. 3x – y = 4 x + 2y = 13 A.(5, 11) B.(–1, 7) C.(3, 5) D.(4, 8)

pooja195 (pooja195):

Hi there! Welcome to openstduy :) So we have 2 equations 3x – y = 4 <---equation 1 x + 2y = 13 <----equation 2 I know the substitution method well enough to teach it to you so lets stick with that :)) Lets start by isolating a variable equation 2 has a variable alone lets isolate it x + 2y = 13 <----equation 2 Subtract 2y from both sides x=-2y+13 Now we have 3x – y = 4 <---equation 1 x=-2y+13 <---equation 2 Now we know the value of x let's plug equation #2 into equation #1 \[\huge~\rm~\bf 3-2y+13 – y = 4 \]Solve for y tell me what you get :)

pooja195 (pooja195):

@jojoheart if you are confused on any part please let me know!

OpenStudy (anonymous):

So is it D?

pooja195 (pooja195):

Nope

OpenStudy (anonymous):

B..?

pooja195 (pooja195):

Nope how bout you try to do the work? :)

OpenStudy (anonymous):

Thanks I got it! It's A!

pooja195 (pooja195):

Its not A -.-